OpenStudy (anonymous):

Solve for x. 8x - 4 = 2x - 3

OpenStudy (anonymous):

subtract \(2x\) from both sides add \(4\) to both sides divide both sides by \(6\)

OpenStudy (anonymous):

okay and how woukd i fugure out the answer

OpenStudy (anonymous):

If you follow satellites instructions you'll get the answer. Look at it like two problems, the first is on the right side of the equal sign and the second on the left, whatever you do to one you have to do to the other. So start out by taking 2x away from both sides.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

Alright, then what do you have?

OpenStudy (anonymous):

umm okay idk if i did this right but 8x - 4 = 2x - 3 -2x -2x ---------------- 6 0 4 4 - - 10 4 /6 /6

OpenStudy (anonymous):

Close! So when you subtract 2x on both sides it leaves you with 6x-4=-3 Which you were definitely on track to get but be careful not to drop the variable too early. The goal once you have x on one side is to get it alone, so after that we add four to both sides which gives us 6x=-3+4 or 6x=1. then divide both sides by 6 to get your answer.

OpenStudy (anonymous):

If you feel like maybe you don't have the right answer these types of problems are nice because you can check them. Just input whatever you got for x on both sides and see if they add up to the same thing. Make sense?

OpenStudy (anonymous):

might 1 be the answer?

OpenStudy (anonymous):

It miiight... but it's not. If you input 1 where x is in both the equations it'll give you different answers for both sides, you want the same number on both sides. The answer is a fraction. so did you see how we got to 6x=1?

OpenStudy (anonymous):

1/6 or 7/6

OpenStudy (anonymous):

1/6. Because when you divide 6x by 6 and 1 by 6, the one can't divide, so you get x=1/6

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Yep!